OpenStudy (anonymous):

simplify 2x^2+20x+32/x^2-2x-80

OpenStudy (anonymous):

\[\frac{ 2(x^2+10x+16) }{x^2-2x-80 }\]

OpenStudy (anonymous):

1*16=16 8+2=10 8*2=16 make factors 1*-80=-80 10-8=2 10*-8=-80 make factors and cancel the common.